Find three consecutive even integers such that when the first integer is multiplied by the third integer, the result is 2 more than 5 times the second integer. (only in algebraic solution)
Answer by stanbon(75887)

Find three consecutive even integers such that when the first integer is multiplied by the third integer, the result is 2 more than 5 times the second integer.
---
1st: 2x-2
2nd: 2x
3rd: 2x+2
-----------
Equation:
(2x-2)(2x+2) = 5(2x)+2
------
4x^2-4 = 10x + 2
4x^2 - 10x -6 = 0
---
2x^2 -5x -3 = 0
(2x+1)(x-3) = 0
---
Positive solution:
x = 3
---
1st: 2x-2 = 4
2nd: 6
3rd: 8
==================
Cheers,
Stan H.
===================